OpenStudy (maheshmeghwal9):

If \[\LARGE{\color{green}{z+\sqrt{2}\space \space |z+1|+i=0.}}\]Then z = ?????

OpenStudy (maheshmeghwal9):

Where "z" is a complex number.

OpenStudy (anonymous):

sqrt{2}|z+1| = -z - i |z+1| is purely real. implying -z-i must be real as well. you can solve the rest now.

OpenStudy (maheshmeghwal9):

didn't gt :(

OpenStudy (anonymous):

take z = x + iy and solve.

OpenStudy (maheshmeghwal9):

yeah i did that

OpenStudy (maheshmeghwal9):

but didn't gt answer

OpenStudy (maheshmeghwal9):

i gt 2 equations at last: - \[y+1=0\]&\[x+\sqrt{2[(x+1)^2+(y)^2]}\]

OpenStudy (maheshmeghwal9):

but didn't gt answer:(

OpenStudy (maheshmeghwal9):

second equation =0*

OpenStudy (klimenkov):

\(x+iy+\sqrt{2}\cdot\sqrt{(x+1)^2+y^2}+i=0\) That means you have a system of 2 equations: \(x+\sqrt{2}\cdot\sqrt{(x+1)^2+y^2}=0\) \(y+1=0\) Solve it!

OpenStudy (maheshmeghwal9):

i have gt the same equation above but didn't gt answer at last:/

OpenStudy (maheshmeghwal9):

u plz tell me wt answer u get?

OpenStudy (maheshmeghwal9):

for x & y?

OpenStudy (klimenkov):

\(z=-2-i\)

OpenStudy (maheshmeghwal9):

oh yeah it is the answer but i m nt getting i put y = -1 in first equation. but i m nt getting x =-2

OpenStudy (klimenkov):

Please, write your detailed solution. By steps.

OpenStudy (maheshmeghwal9):

ok! oops sorry i gt answer while writing steps actually big mistake of plus / minus by me sorry! thanx for the help:)